LadyRavyRay:

Find the volume of the rectangular prism. Express your answer as a simplified mixed number. 8 3/8m x 4m x 8 1/4 m V = _____ cubic meters

snowflake0531:

Do you want me to just check this for you or..?

AZ:

Just multiply the three numbers ta-da

LadyRavyRay:

I don't understand how to multiple that.

snowflake0531:

3/8 is also 0.375 *shrug Or you could multiply it in fractions, but that's annoying

LadyRavyRay:

Decimals are better.

snowflake0531:

Then it's just 8.375 * 4 * 8.25

AZ:

you can use a calculator 3/8 = 0.375 so 8 3/8 = 8.375 and 1/4 = 0.25 so 8 1/4 = 8.25

LadyRavyRay:

Is the answer, 276.375

LadyRavyRay:

But I have to make that a mixed number right?

snowflake0531:

0.375 = 3/8

LadyRavyRay:

276 3/8?

AZ:

yes

snowflake0531:

ye